# The Larger the Better? Improved LLM Code-Generation via Budget Reallocation

### v1

@@ 7 Discussion & Limitations
Our results show that using smaller models with the same amount of compute can improve LLM code-generation performance. An interesting question we do not fully address is whether, given enough compute, the larger models will overtake the smaller ones, or perhaps they will all saturate at a similar performance level at some point. Our HumanEval and MBPP results seem to slightly support the latter hypothesis (as all models begin to saturate, see Figures 2 and 3). However, unfortunately, due to compute constraints, our setting is restricted to exploring only a limited number of generations per model.222For instance, generating answers for the examples of the APPS benchmark with a B model takes about days using a node of A GPUs. We note that despite this limitation, in practice, due to these costs our conclusions apply to most practical use-cases. We defer more expensive experiments to future work.
@@ 8 Conclusion
In this work, we compared large language models with smaller-sized models under fixed budget constraints (i.e., FLOPs and wall-time). We evaluated the models using execution-based code-generation tasks, which provide access to unit-tests. Our findings reveal that generating multiple outputs from a B model may lead to gains of up to 15% over a single generation from a B model across five tasks. This highlights the potential of using smaller models instead of larger ones. In scenarios where unit tests or other solution verifiers are unavailable, we explored a simple ranking-based approach for candidate selection. We found the proposed ranking approach falls short in performance compared to a single output from the larger model. Our findings emphasize the importance of studying approaches for ranking LLM outputs, which hold great potential to not only improve model performance but also improve budget allocation. To further enhance this research direction we release over M samples from the Code Llama 7B models considering both HumanEval and MBPP benchmarks.

### v2

@@ 7 Discussion & Limitations
Our results show that using smaller models with the same amount of compute can improve LLM code-generation performance. An interesting question we do not fully address is whether, given enough compute, the larger models will overtake the smaller ones, or perhaps they will all saturate at a similar performance level at some point. Our HumanEval and MBPP results seem to slightly support the latter hypothesis (as all models begin to saturate, see Figures 2 and 3). However, unfortunately, due to compute constraints, our setting is restricted to exploring only a limited number of generations per model.666For instance, generating answers for the examples of the APPS benchmark with a B model takes about days using a node of A GPUs. We note that despite this limitation, in practice, due to these costs our conclusions apply to most practical use-cases. We defer more expensive experiments to future work.
@@ 8 Conclusion
In this work, we compared large language models with smaller-sized models under fixed budget constraints (i.e., FLOPs and wall-time). We evaluated the models using execution-based code-generation tasks, which provide access to unit-tests. Our findings reveal that generating multiple outputs from a B model may lead to gains of up to 15% over a single generation from a B model across five tasks. This highlights the potential of using smaller models instead of larger ones. In scenarios where unit tests or other solution verifiers are unavailable, we explored a simple ranking-based approach for candidate selection. We found the proposed ranking approach falls short in performance compared to a single output from the larger model. Our findings emphasize the importance of studying approaches for ranking LLM outputs, which hold great potential to not only improve model performance but also improve budget allocation. To further enhance this research direction we release over M samples from the Code Llama 7B models spanning both HumanEval and MBPP benchmarks.
9 Acknowledgments
We thank Miri Varshavsky Hassid for the great feedback and moral support.
Appendix A Llama-3 Results
We present Llama-3 results for the HumanEval and MBPP benchmarks in Figures 10 and 11, respectively.
Appendix B Detailed pass@ Results
In Tables 2, 3 and 4 presents precise pass@ results for the datasets examined (HumanEval, MBPP and APPS, respectively). Due to the infeasibility of reporting results for all , we provide results for selected values. Nevertheless, it is important to note that all relevant values were calculated and used in the computation of the figures.
1 | 2 | 4 | 16 | 64 | 128 | 256 | 500 | 1000 | |
Code Llama B | 28.2 | 38.5 | 48.9 | 68.7 | 83.9 | 89.0 | 92.7 | 95.0 | 96.3 |
Code Llama B | 32.7 | 44.4 | 56.2 | 77.0 | 89.0 | 92.3 | 94.8 | 96.5 | –.- |
Code Llama B | 38.6 | 51.4 | 63.4 | 81.8 | 91.7 | 94.3 | –.- | –.- | –.- |
Code Llama B | 46.7 | 61.1 | 73.2 | 87.3 | 94.4 | –.- | –.- | –.- | –.- |
Llama-3 B | 31.8 | 43.1 | 54.4 | 74.0 | 86.9 | 90.9 | 94.0 | 96.0 | 97.3 |
Llama-3 B | 47.3 | 59.3 | 69.8 | 84.7 | 92.8 | –.- | –.- | –.- | –.- |
1 | 2 | 4 | 16 | 64 | 128 | 256 | 500 | 1000 | |
Code Llama B | 37.3 | 48.0 | 57.3 | 71.6 | 81.1 | 84.8 | 87.9 | 90.3 | 92.4 |
Code Llama B | 42.3 | 53.4 | 62.5 | 75.3 | 84.0 | 87.4 | 90.4 | 92.8 | –.- |
Code Llama B | 49.2 | 60.1 | 68.6 | 79.2 | 85.8 | 88.7 | –.- | –.- | –.- |
Code Llama B | 57.3 | 67.7 | 74.8 | 83.7 | 90.0 | –.- | –.- | –.- | –.- |
Llama-3 B | 41.1 | 52.4 | 61.8 | 75.5 | 84.4 | 87.7 | 90.5 | 92.8 | 94.7 |
Llama-3 B | 61.2 | 71.5 | 78.7 | 87.2 | 92.9 | –.- | –.- | –.- | –.- |
1 | 2 | 4 | 16 | 32 | 64 | 128 | 256 | 500 | |
APPS-introductory | |||||||||
Code Llama B | 4.5 | 7.7 | 12.1 | 23.0 | 28.6 | 34.0 | 39.3 | 44.4 | 49.0 |
Code Llama B | 10.3 | 16.3 | 23.1 | 36.3 | 42.3 | 47.8 | 52.8 | –.- | –.- |
Code Llama B | 12.6 | 19.7 | 27.4 | 40.8 | 46.8 | 52.5 | –.- | –.- | –.- |
Code Llama B | 22.6 | 30.5 | 38.1 | 52.5 | 58.5 | –.- | –.- | –.- | –.- |
APPS-interview | |||||||||
Code Llama B | 0.9 | 1.6 | 2.7 | 6.1 | 8.3 | 10.8 | 13.6 | 16.4 | 19.2 |
Code Llama B | 2.2 | 3.8 | 5.9 | 11.6 | 14.8 | 18.3 | 21.9 | –.- | –.- |
Code Llama B | 2.9 | 4.9 | 7.6 | 14.4 | 18.1 | 22.0 | –.- | –.- | –.- |
Code Llama B | 5.5 | 8.7 | 12.6 | 21.8 | 26.6 | –.- | –.- | –.- | –.- |
APPS-competition | |||||||||
Code Llama B | 0.1 | 0.3 | 0.5 | 1.6 | 2.6 | 4.1 | 6.1 | 8.6 | 11.6 |
Code Llama B | 0.6 | 1.2 | 2.1 | 5.3 | 7.8 | 10.9 | 14.4 | –.- | –.- |
Code Llama B | 0.7 | 1.3 | 2.3 | 6.1 | 9.0 | 12.6 | –.- | –.- | –.- |
Code Llama B | 1.7 | 3.0 | 4.9 | 11.1 | 15.3 | –.- | –.- | –.- | –.- |